WRITE AS A COMPLEX NUMBER IN STANDARD FORM ;)
(5+3i)^2

Answer by Alan3354(69443)

WRITE AS A COMPLEX NUMBER IN STANDARD FORM ;)
(5+3i)^2
-----------
= (5+3i)*(5+3i)
= 5*(5+3i) + 3i*(5+3i)
= 25+15i + 15i-9
= 16 + 30i